OpenStudy (anonymous):

write an equation of the line that is perpendicular to 6x-5y=7 that passes through the point (-11,5) using point slope form

OpenStudy (texaschic101):

you need to find the slope first....do you know how to do that ?

OpenStudy (texaschic101):

do as Hero said and put 6x - 5y = 7 into slope intercept form which is y = mx + b. The position of m is the slope. Can you put it into that form ?

OpenStudy (anonymous):

slope = y2-y1/x2-x1 right

OpenStudy (texaschic101):

that is the slope formula, but we do not have 2 points, therefore, we need that given equation in y = mx + b form

OpenStudy (anonymous):

ok cam you show me so i dont make any mistakes. i just want to be sure so i understand

OpenStudy (texaschic101):

6x - 5y = 7 (subtract 6x from both sides) -5y = -6x + 7 (now divide by -5 to get y by itself) y = -6/-5x - 7/5 y = 6/5x - 7/5 do you understand how I did that ?

OpenStudy (anonymous):

yes i do

OpenStudy (texaschic101):

y = mx + b form and m is your slope, therefore, the slope of this equation is 6/5. However, we are looking for a perpendicular line so the slope we need has to be the negative reciprocal of the slope of the equation. All that means is " flip " the slope and change the sign. slope is 6/5 and the slope we need is -5/6. Understand ?

OpenStudy (anonymous):

Yes, now that is how i do the slope form right? Now what about the coordinates and the point slope form?

OpenStudy (texaschic101):

point slope form is : y - y1 = m(x - x1) slope(m) = -5/6 (-11,5) x1 = -11 and y1 = 5 now sub your info in the formula and what do you get ? You do not have to solve it, just put it in the formula, parenthesis and all.

OpenStudy (texaschic101):

are you stuck ?

OpenStudy (anonymous):

y-5= 6/5(X-(-11)

OpenStudy (texaschic101):

not quite....we are using the negative reciprocal of the slope 6/5, which is -5/6. So the point slope formula is : y - 5 = -5/6(x -(-11) or you could probably just put : y - 5 = -5/6(x + 11)

OpenStudy (anonymous):

Oh ok thank you so much, now i dont want to be a bother but do you mind helping me with other questions :)

OpenStudy (texaschic101):

possibly...hope they are not geometry related...not good at those

OpenStudy (anonymous):

i will post it and we will see i can. Thank you

OpenStudy (texaschic101):

no problem :)